OpenStudy (donnie1999):

-------- WILL MEDAL -------- -------- HELP PLEASE -------- Find the x-intercepts for the parabola defined by the equation below. y = x^2 - 9x + 18

OpenStudy (donnie1999):

@butterflydreamer

OpenStudy (anonymous):

What do you think? What did you get after working it out?

OpenStudy (anonymous):

@UsukiDoll

OpenStudy (anonymous):

First ask your self what two factors when you times them equals 18 and also sums(add together) up to -9

rishavraj (rishavraj):

to find x-intercept ...let y = 0

rishavraj (rishavraj):

so solve x^2 - 9x + 18 = 0

OpenStudy (donnie1999):

i got (0,3) and (0,6)

rishavraj (rishavraj):

dont u think it should be (3 , 0) and (6 , 0)???

OpenStudy (usukidoll):

to find x - intercept let the y = 0 (x^2-9x+18) = 0 so the y's have to be 0 you are solving for the x values

OpenStudy (usukidoll):

consider the 2 numbers 6 and 3. Those numbers work because 6 x 3 =18 (got my last term) 6+3 =9 (got my middle term) so now we can factor (x-6)(x-3) =0 so now we have 2 cases x - 6 =0 and x -3 =0 solve for x and we should have what @rishavraj said

OpenStudy (usukidoll):

the y is fixed at 0 for these cases... since we are looking for x-intercept